OpenStudy (anonymous):

The Height , H Of A Baksetball , in meters at t seconds after it is tossed out of a window is modeled by : H=-5t^2 + 20t +15. A Boy Shoots At The Baseball With A Paintball Gun . The Trajectory Of The Paintball Is Given By The Equation : H=3t + 3 . Will The Paintball Hit The Baseball ? IF So When? At What Height Will The Baseball bE ? @jtvatsim

OpenStudy (jtvatsim):

Alrighty, what have we got here...

OpenStudy (jtvatsim):

"something is flying through the air, somebody does something to hit in midair" yep same problem type... :)

OpenStudy (anonymous):

Yup Sooo... -5t^2 +20t+15=3t +3

OpenStudy (jtvatsim):

Yessiree! Will the heights be equal? And when? Let's find out. The audience waits in suspense...

OpenStudy (anonymous):

Lets Fine Outt Next Step Add ?

OpenStudy (jtvatsim):

Either add or subtract, whatever will make the other side 0. Try something, and I'll let you know what I think. :)

OpenStudy (anonymous):

Ok Subtract Because there is only one negative .

OpenStudy (jtvatsim):

You are right to subtract. But not because there is only one negative. Let's pause on this step to make sure you've got this part down. That way, you'll always have the confidence to get through this stage. :)

OpenStudy (jtvatsim):

When I see -5t^2 + 20t + 15 = 3t + 3, I immediately focus my attention on the right side of the equation (look at 3t + 3).

OpenStudy (anonymous):

Ok

OpenStudy (jtvatsim):

Then, ask yourself, I see a 3t, how can I get rid of it? In this case I could just subtract a 3t from both sides like this: -5t^2 + 20t + 15 = 3t + 3 -3t -3t ___________________________ -5t^2 + 17t + 15 = 3 Then my 3t is gone!

OpenStudy (jtvatsim):

But, then, of course, I still see a pesky 3 there. How could I get rid of that since I just want zero on that side? Well... subtract a 3 from both sides, just like we did before! -5t^2 + 17t + 15 = 3 -3 -3 -------------------- -5t^2 + 17t + 12 = 0

OpenStudy (anonymous):

Ohhh i c

OpenStudy (jtvatsim):

Great! It's not too hard once you see the pattern. But it is a little hard to say in words. :)

OpenStudy (jtvatsim):

Here's one more example just to make sure you've got it. -66t^2 + 24t + 31 = -6t + 4 +6t +6t --------------------------- -66t^2 + 30t + 31 = 4 -4 -4 ---------------------- -66t^2 + 30t + 27 = 0

OpenStudy (jtvatsim):

In this case, I added to get rid of the -6t, but subtracted to get rid of the +4. Any questions on that process?

OpenStudy (jtvatsim):

And of course, what I just wrote above was a made-up example and has nothing to do with "our" current problem. :)

OpenStudy (anonymous):

i noticed Lol Those Off #;s

OpenStudy (jtvatsim):

Haha! I tried to make it obvious that I was going crazy there. ;) Anyways, we've got our equation now: -5t^2 + 17t + 12 = 0

OpenStudy (jtvatsim):

From what I've seen, that should be the end of the "hard" part mathematically. The rest of the problem is just making sure our calculator does the right thing. :)

OpenStudy (anonymous):

Yup Thats The Last Thing 2 do

OpenStudy (jtvatsim):

OK, then. Once you have your proposed solutions from the quadratic formula, we'll compare notes.

OpenStudy (anonymous):

Done

OpenStudy (jtvatsim):

Awesome! What did you get?

OpenStudy (anonymous):

x1=-0.5 x2 :4

OpenStudy (jtvatsim):

Yep, that's pretty close to what I got, x1 = -0.59999 x2 = 4

OpenStudy (anonymous):

i got the 9999 2 lol

OpenStudy (anonymous):

didnt think i needed it

OpenStudy (jtvatsim):

Oh, well, look at you! It's slightly different rounding -0.59 rounds to -0.6 (then I wouldn't have said anything...) You can feel that I'm becoming more nit picky! :) That's a good thing, it means you are moving beyond the basics. :)

OpenStudy (anonymous):

Lol Im Trying .. Another One Then Ill Leave You Alone Lol

OpenStudy (jtvatsim):

No worries! Glad to help. It's fun watching someone start out struggling, but then learn and start mastering it. It's really cool to see happen. :)

OpenStudy (jtvatsim):

So, we've got our possible solutions, but we need to check them. Here goes...

OpenStudy (jtvatsim):

Looks like t = -0.6 gives h = 1.2 and t = 4 gives h = 15

OpenStudy (anonymous):

The Correct One Is ....4

OpenStudy (jtvatsim):

I'd agree, even though the heights are both positive (and therefore possible), the negative time is weird. The answer is Yes, time is 4, and height is 15. Yahoo! :D

OpenStudy (anonymous):

Yay !!!!!!!!!! You Helped Me Sooo Much ..After This Next One Ill Have Just Four More ...Wouldve Been Up All Night

OpenStudy (jtvatsim):

You are welcome! Based on looking at the problems are they relatively similar? If not, I could help you look at one that is a different type if you feel comfortable with these "objects flying and hitting" questions.

OpenStudy (anonymous):

This Next One Is A Different One .

OpenStudy (anonymous):

Is a summer packet soo

OpenStudy (jtvatsim):

Let's take a look at it then if you want. Even if I have to log off soon, I can at least try to come up with a plan of attack. :)

OpenStudy (anonymous):

Sure

OpenStudy (jtvatsim):

sounds good